# A Configuration-First Framework for Reproducible, Low-Code Localization

### v1

@@ 6.1. Cross–platform qualitative comparison
R1: Low-code operation Barrier to entry for running and iterating on end-to-end localization experiments via configuration (or GUI) rather than code; minimal setup and boilerplate. ”Very high” = little/no coding; ”Very low” = substantial coding.
R2: Reproducibility by default Ability to reproduce identical results by default: versioned code/config/data, captured environments, controlled randomness, isolated execution, and automatic artifact/metrics recording. ”Very high” = full reproducibility guaranteed without extra effort; ”Very low”= reproducibility depends entirely on manual intervention.
R3: Common evaluation layer Built-in, reusable evaluation across methods and datasets (metrics, reporting, artifact schema) without bespoke glue code. ”Very high” = rich, standardized evaluation layer that works across models/datasets out of the box; ”Very low” = no shared evaluation.
R4: Built-in extensibility Effort to add or replace models, metrics, datasets, or stages through configuration or small adapters without modifying core internals. ”Very high” = seamless extensibility via lightweight configuration; ”Very low” = extensions require invasive modifications to the framework’s internals.
R5: Fast setup and iteration The effort required for initial setup and to iterate on experiments. ”Very high” = ready-to-use environment with minimal dependencies; ”Very low” = heavy installation burden, long configuration times, and slow iteration cycles.
We evaluate each platform on five requirement-oriented criteria (R1–R5) plus a descriptive Primary focus, using the definitions in Table 1. To facilitate a fair cross-platform assessment (including GUI-first tools), the wording in Table 1 is broadened relative to Section 3. For each item, we assign an ordinal grade from ”Very low” to ”Very high”, based on official documentation and hands-on use. When a capability is not available out of the box but can be reasonably achieved via supported extensions or common external tools, the grade is capped at ”Medium”. This comparison is intended to situate LOCALIZE among existing solutions and to assess, at a high level, how well each platform aligns with the requirements that motivate our framework.
6.2. Materials
All the quantitative experiments were carried out partly on the UMU cellular localization (Bertalanič et al., 2022) dataset and primarily on the LOG-a-TEC Bluetooh localization dataset (Asensio-Garriga et al., 2025).
6.3. Lines of code study
We implemented functionally equivalent baseline pipelines in both environments. Both produced identical outputs and were initially composed of the stages Prepare, Featurize, Split, and Grid-search for the UMU dataset, implemented as they were described in Section 5. The pipelines initially used the LinearRegression model from scikit-learn, a two-value hyperparameter grid for fit_intercept, evaluated with 5-fold KFold and rmse. We then introduced four incremental changes:
  • Change 1: replace the estimator and add a second one,
  • Change 2: switch the dataset from UMU to LOG-a-TEC,
6.4. Runtime and resource profiling
6.5. Scalability on LOG-a-TEC
[…]
6.6. Reproducibility measures
7. Evaluation
In this section, we discuss the evaluation result of the LOCALIZE framework introduced in Section 4 according to the methodology elaborated in Section 6. First, we focus on the qualitative cross–platform comparison of five solutions, followed by a quantitative study against a Jupyter Notebook baseline, and concluding with a benchmark on the LOG-a-TEC dataset scaled to different sizes.
7.1. Cross–platform qualitative comparison
We evaluate five platforms, LOCALIZE, WEKA, Orange3, Jupyter notebooks, and Ludwig, against the five requirements (R1–R5) defined in Table 1, using the rubric in Section 6.1. Table 2 lists platforms across columns and requirements down the rows. Below, we explain the grades and the main takeaways for each requirement.
Table 2. Qualitative comparison of platforms along evaluation features (definitions in Table 1).
Feature/Platform LOCALIZE WEKA (Witten et al., 2025) Orange3 (Demsˇar et al., 2013) Jupyter notebook (Granger and Pérez, 2021) Ludwig (Molino et al., 2019)
@@ 7.1. Cross–platform qualitative comparison
As per first row in Table 2, LOCALIZE targets end-to-end localization workflows. WEKA emphasizes general classical ML, Orange3 focuses on visual data mining, Jupyter emphasizes code-centric experimentation, and Ludwig targets deep learning with configuration files. This establishes scope and is not scored.
@@
As can be seen in second row of Table 2, both WEKA and Orange3 have a ”Very high” score for R1, as users can assemble pipelines in a GUI with minimal coding. Jupyter has a score of ”Very low” since routine edits require writing and maintaining scripts. LOCALIZE and Ludwig achieve a score of ”High”, since both support configuration-first editing for common changes while allowing custom code when needed.
@@
Analysing the third row of Table 2, it can be seen that LOCALIZE has a ”High” score due to configuration, environments, execution isolation, and artifacts being handled as part of the default workflow. In comparison, WEKA score is ”Low”, as it only provides limited reproducibility support, because GUI projects are not inherently versioned, and environment capture is manual. Orange3 sits in the middle at score ”Medium”, as workflows can be saved, but data and environment versioning require external practice. Jupyter notebooks are rated as ”Low” due to offering no built-in reproducibility support, so they depend on user discipline and external tooling for reproducibility. However, even with extensions, achieving end-to-end replicability typically requires a multi-step setup and careful consideration. Finally, Ludwig’s YAML promotes consistent experiment specs, but environment and dataset versioning remain external concerns, so it achieves score rate of ”Medium”.
@@
LOCALIZE scores ”Very high” due to uniform metrics, reporting, and artifact schema applied across datasets and methods. Orange3 and Ludwig are ”Medium”. Orange3 provides evaluation widgets, yet consistent reuse across heterogeneous methods and datasets often needs add-ons or custom wiring. Ludwig supplies a strong evaluation for its deep-learning workflows, but does not generalize to classical pipelines out of the box. WEKA and Jupyter are ”Low” because comparable, reusable evaluation generally requires bespoke code or plug-ins. The takeaway is that LOCALIZE uniquely enforces like-for-like scoring and reports without extra glue.
@@
Jupyter and Ludwig are ”Very high” since they are designed for heavy customization. LOCALIZE also scores ”Very high” by exposing clear extension points that can be referenced from configuration, so additions plug into the same orchestration and reporting. WEKA is ”High” because it is extensible through Java APIs and plug-ins, although GUI integration can be involved. Orange3 is ”Medium” due to the overhead of developing and maintaining PyQt widgets.
@@
LOCALIZE is rated ”Very high”, as it combines ready-to-run datasets and templates with cached, stage-wise execution, so common edits avoid full reruns. Jupyter and Ludwig iterate quickly once code patterns and environments are in place and are therefore rated ”High”. WEKA is rate ”High” due to GUIs support of quick initial setup, although complex studies add manual steps. Orange3 is ”Medium” since visual assembly scales less smoothly when experiments branch or require custom tooling.
Table 2 highlights a consistent trade-off. GUI-first tools make initial use easy but rely on user practice for exact repeatability and uniform scoring. Code-first tools enable deep customization and rapid iteration, but place reproducibility and evaluation consistency on the user. LOCALIZE is designed to close these gaps by making reproducibility the default, while preserving low-code operation and extensibility. In the context of localization studies, this combination makes LOCALIZE the most appropriate choice among the compared platforms.
7.2. Comparison against notebooks
@@
Figure 3. Lines of code added (positive) and removed (negative) when introducing successive changes in the Framework and Notebook implementations.:
@@
Figure 3 compares edit effort for four incremental changes in LOCALIZE versus a Jupyter baseline. The x-axis enumerates Changes 1–4; the y-axis reports lines of code, counted as described in Section 6.3. For each platform and change, the upward bar shows lines added and the downward bar shows lines removed, both have their corresponding value for lines of code added/removed written on them. Two bars are plotted per platform, so additions and deletions are visible separately.
@@
Figure 4. Memory usage statistics per pipeline stage for LOCALIZE and the Notebook implementation. Bars show means with first/third quartiles as error bars.:
@@
CPU and wall time by stage.
Figure 5. Wall-clock and CPU time per pipeline stage for LOCALIZE and the Notebook implementation.:
7.3. Scalability on LOG-a-TEC
@@
8. Limitations and Future Work

### v2

@@ V-E Preconfigured datasets
R1: Low-code operation Barrier to entry for running and iterating on end-to-end localization experiments via configuration (or GUI) rather than code; minimal setup and boilerplate. "Very high" = little/no coding; "Very low" = substantial coding.
R2: Reproducibility by default Ability to reproduce identical results by default: versioned code/config/data, captured environments, controlled randomness, isolated execution, and automatic artifact/metrics recording. "Very high" = full reproducibility guaranteed without extra effort; "Very low"= reproducibility depends entirely on manual intervention.
R3: Common evaluation layer Built-in, reusable evaluation across methods and datasets (metrics, reporting, artifact schema) without bespoke glue code. "Very high" = rich, standardized evaluation layer that works across models/datasets out of the box; "Very low" = no shared evaluation.
R4: Built-in extensibility Effort to add or replace models, metrics, datasets, or stages through configuration or small adapters without modifying core internals. "Very high" = seamless extensibility via lightweight configuration; "Very low" = extensions require invasive modifications to the framework’s internals.
R5: Fast setup and iteration The effort required for initial setup and to iterate on experiments. "Very high" = ready-to-use environment with minimal dependencies; "Very low" = heavy installation burden, long configuration times, and slow iteration cycles.
TABLE II: Qualitative comparison of platforms along evaluation features (definitions in Table I).
Feature/Platform LOCALIZE WEKA [weka_book] Orange3 [orange_pdf] Jupyter notebook [granger_jupyter_2021] Ludwig [ludwig_pdf]
@@
VI Evaluation methodology
In this section we elaborate on the methodology to evaluate the proposed LOCALIZE framework introduced in Section IV along the following four aspects: (i) cross-tool capabilities using a qualitative comparison, (ii) developer effort by comparing our framework with a plain Jupyter notebook under controlled edits, (iii) runtime and resource usage by pipeline stage, and (iv) scalability as dataset size increases.
VI-A Cross–platform qualitative comparison
As the first step in the evaluation of LOCALIZE, we identify the closest platforms that are strong candidates to enable similar functionality as the proposed framework and dive deeper to understand specific differences. We identify WEKA [weka_book], Orange3 [orange_pdf], Ludwig [ludwig_pdf], and Jupyter notebooks in JupyterLab [granger_jupyter_2021] as the closest relevant platforms.
We evaluate each platform on five requirement-oriented criteria (R1–R5) plus a descriptive Primary focus, using the definitions in Table I. To facilitate a fair cross-platform assessment (including GUI-first tools), the wording in Table I is broadened relative to Section III. For each item, we assign an ordinal grade from "Very low" to "Very high", based on official documentation and hands-on use. When a capability is not available out of the box but can be reasonably achieved via supported extensions or common external tools, the grade is capped at "Medium". This comparison is intended to situate LOCALIZE among existing solutions and to assess, at a high level, how well each platform aligns with the requirements that motivate our framework.
VI-B Materials
All the quantitative experiments were carried out partly on the UMU cellular localization [morano_outdoor_2022] dataset and primarily on the LOG-a-TEC Bluetooh localization dataset [asensio-garriga_university_2025].
VI-C Lines of code study
We implemented functionally equivalent baseline pipelines in both environments. Both produced identical outputs and were initially composed of the stages Prepare, Featurize, Split, and Grid-search for the UMU dataset, implemented as they were described in Section V. The pipelines initially used the LinearRegression model from scikit-learn, a two-value hyperparameter grid for fit_intercept, evaluated with 5-fold KFold and rmse. We then introduced four incremental changes:
  • Change 1: replace the estimator and add a second one,
  • Change 2: switch the dataset from UMU to LOG-a-TEC,
VI-D Runtime and resource profiling
VI-E Scalability on LOG-a-TEC
[…]
VI-F Reproducibility measures
VII Evaluation
In this section, we discuss the evaluation result of the LOCALIZE framework introduced in Section IV according to the methodology elaborated in Section VI. First, we focus on the qualitative cross–platform comparison of five solutions, followed by a quantitative study against a Jupyter Notebook baseline, and concluding with a benchmark on the LOG-a-TEC dataset scaled to different sizes.
VII-A Cross–platform qualitative comparison
We evaluate five platforms, LOCALIZE, WEKA, Orange3, Jupyter notebooks, and Ludwig, against the five requirements (R1–R5) defined in Table I, using the rubric in Section VI-A. Table II lists platforms across columns and requirements down the rows. Below, we explain the grades and the main takeaways for each requirement.
@@ VII-A Cross–platform qualitative comparison
As per first row in Table II, LOCALIZE targets end-to-end localization workflows. WEKA emphasizes general classical ML, Orange3 focuses on visual data mining, Jupyter emphasizes code-centric experimentation, and Ludwig targets deep learning with configuration files. This establishes scope and is not scored.
@@
As can be seen in second row of Table II, both WEKA and Orange3 have a "Very high" score for R1, as users can assemble pipelines in a GUI with minimal coding. Jupyter has a score of "Very low" since routine edits require writing and maintaining scripts. LOCALIZE and Ludwig achieve a score of "High", since both support configuration-first editing for common changes while allowing custom code when needed.
@@
Analysing the third row of Table II, it can be seen that LOCALIZE has a "High" score due to configuration, environments, execution isolation, and artifacts being handled as part of the default workflow. In comparison, WEKA score is "Low", as it only provides limited reproducibility support, because GUI projects are not inherently versioned, and environment capture is manual. Orange3 sits in the middle at score "Medium", as workflows can be saved, but data and environment versioning require external practice. Jupyter notebooks are rated as "Low" due to offering no built-in reproducibility support, so they depend on user discipline and external tooling for reproducibility. However, even with extensions, achieving end-to-end replicability typically requires a multi-step setup and careful consideration. Finally, Ludwig’s YAML promotes consistent experiment specs, but environment and dataset versioning remain external concerns, so it achieves score rate of "Medium".
@@
LOCALIZE scores "Very high" due to uniform metrics, reporting, and artifact schema applied across datasets and methods. Orange3 and Ludwig are "Medium". Orange3 provides evaluation widgets, yet consistent reuse across heterogeneous methods and datasets often needs add-ons or custom wiring. Ludwig supplies a strong evaluation for its deep-learning workflows, but does not generalize to classical pipelines out of the box. WEKA and Jupyter are "Low" because comparable, reusable evaluation generally requires bespoke code or plug-ins. The takeaway is that LOCALIZE uniquely enforces like-for-like scoring and reports without extra glue.
@@
Jupyter and Ludwig are "Very high" since they are designed for heavy customization. LOCALIZE also scores "Very high" by exposing clear extension points that can be referenced from configuration, so additions plug into the same orchestration and reporting. WEKA is "High" because it is extensible through Java APIs and plug-ins, although GUI integration can be involved. Orange3 is "Medium" due to the overhead of developing and maintaining PyQt widgets.
@@
LOCALIZE is rated "Very high", as it combines ready-to-run datasets and templates with cached, stage-wise execution, so common edits avoid full reruns. Jupyter and Ludwig iterate quickly once code patterns and environments are in place and are therefore rated "High". WEKA is rate "High" due to GUIs support of quick initial setup, although complex studies add manual steps. Orange3 is "Medium" since visual assembly scales less smoothly when experiments branch or require custom tooling.
Table II highlights a consistent trade-off. GUI-first tools make initial use easy but rely on user practice for exact repeatability and uniform scoring. Code-first tools enable deep customization and rapid iteration, but place reproducibility and evaluation consistency on the user. LOCALIZE is designed to close these gaps by making reproducibility the default, while preserving low-code operation and extensibility. In the context of localization studies, this combination makes LOCALIZE the most appropriate choice among the compared platforms.
VII-B Comparison against notebooks
@@
Figure 3: Lines of code added (positive) and removed (negative) when introducing successive changes in the Framework and Notebook implementations.
Figure 4: Wall-clock and CPU time per pipeline stage for LOCALIZE and the Notebook implementation.
@@
Figure 3 compares edit effort for four incremental changes in LOCALIZE versus a Jupyter baseline. The x-axis enumerates Changes 1–4; the y-axis reports lines of code, counted as described in Section VI-C. For each platform and change, the upward bar shows lines added and the downward bar shows lines removed, both have their corresponding value for lines of code added/removed written on them. Two bars are plotted per platform, so additions and deletions are visible separately.
@@
Figure 5: Memory usage statistics per pipeline stage for LOCALIZE and the Notebook implementation. Bars show means with first/third quartiles as error bars.
[…]
CPU and wall time by stage
@@
VII-C Scalability on LOG-a-TEC
[…]
Per-stage CPU and wall time
@@ VII-C Scalability on LOG-a-TEC
VIII Limitations and Future Work